# Sparse Offline Reinforcement Learning with Corruption Robustness

### 07 Mar 2026 v2

@@ 1.1 Related works
(Ye et al., 2023) $\widetilde{O}\!\left(\tfrac{H^{2}\sqrt{d}\epsilon}{\xi}\right)$ $\mathrm{poly}(d,H)$ Uniform coverage
(Zhang et al., 2021) $\widetilde{O}\!\left(\tfrac{H^{2}\sqrt{d}\epsilon}{\xi}\right)$ $\mathrm{poly}(d,H)$ Uniform coverage
Ours $\widetilde{O}\!\left(\tfrac{H^{2}s\sqrt{\epsilon}}{\xi}\right)$ $\mathrm{poly}(s,H)$ Uniform coverage
(Zhang et al., 2021) $\widetilde{O}\!\left(H^{2}\sqrt{d\kappa\epsilon}\right)$ $\mathrm{poly}(d,H)$ Single-policy concentrability
Ours $\widetilde{O}\!\left(H^{2}\sqrt{s\kappa\epsilon}\right)$ $\mathrm{poly}(s,H)$ Single-policy concentrability
@@
In this paper, we consider data poisoning attacks in offline RL (Zhang et al., 2021). Previous work on corruption-robust offline RL has primarily considered the low-dimensional setting where $N\gg d$ (Zhang et al., 2021; Ye et al., 2023). In contrast, this work focuses on the high-dimensional setting $d>N$. In particular, we employ a value function estimation approach (Zhang et al., 2021) based on sparse robust estimation (Merad and Gaïffas, 2022). We find that, under single-policy concentrability, in sparse settings, integrating sparse oracles into LSVI-type algorithms is considerably more challenging and may yield suboptimal bounds that depend polynomially on $d$, in contrast to prior work in the low-dimensional setting (Zhang et al., 2021). Moreover, we demonstrate that sparse robust estimation can be naturally integrated into the AC framework, effectively eliminating polynomial dependence on $d$. A systematic comparison of our results with previous work on corruption-robust offline RL is provided in Table 1.
@@ 2 Preliminary
We consider the offline setting with contaminated data. Let $\widetilde{\mathcal{D}}$ be a clean dataset consisting of $N$ trajectories, i.e., $\widetilde{\mathcal{D}}=\{(x^{\tau}_{h},a_{h}^{\tau},R_{h}^{\tau})\}_{\tau=1,h=1}^{N,H}$. Suppose there is an adversary who observes the dataset $\widetilde{\mathcal{D}}$ and may arbitrarily corrupt up to $\epsilon N$ trajectories, resulting in a corrupted dataset $\mathcal{D}$. We also denote $\mathcal{D}_{h}$ be data points at horizon $h$. In this work, regarding clean data set, we assume there is an exploratory policy $\pi_{\nu}$, with the occupancy probability $\nu=(\nu_{h})_{h=1}^{H}$. Define $\Sigma_{h}\triangleq\mathbb{E}_{\nu_{h}}\!\left[\phi(x,a)\phi(x,a)^{\top}\right]$ as the covariance matrix of the underlying data distribution $\nu$ at horizon $h$. We introduce an assumption on the distribution $\nu$, which is crucial for our main result.
@@ 3.3 Using SRLE Estimators in Offline Reinforcement Learning
$\mathcal{D}^{\pi}_{h}\triangleq\!\left\{\!\left(\underbrace{\phi(x_{h}^{\tau},a_{h}^{\tau})}_{z_{i}},\;\\ \underbrace{R_{h}^{\tau}+\mathbb{E}_{\pi_{h+1}}\!\left[F(x^{\tau}_{h+1},a)\right]}_{y_{i}}\right)\right\}_{\tau=1}^{N}.$
@@
$\mathcal{D}^{*}_{h}$ $\triangleq\!\left\{\!\left(\phi(x_{h}^{\tau},a_{h}^{\tau}),\;R_{h}^{\tau}+\max_{a\in\mathcal{A}}F(x^{\tau}_{h+1},a)\right)\right\}_{\tau=1}^{N}.$ (6)
@@ 4.1 Sparse Robust LSVI with Uniform Coverage
$\mathrm{SubOpt}(\pi_{\star},\widehat{\pi})=O\left(\frac{H^{2}s\log(d/\delta)}{\xi\sqrt{N}}+\frac{H^{2}s\sqrt{\epsilon}}{\xi}\right).$ (7)
@@ 5.1 Sparse Robust AC with Uniform Coverage
Suppose Assumption 2.1 and 2.2 hold. Then, with step size $\eta=\sqrt{\frac{\log\mathcal{A}}{N}}$, after $T=N$ iterations the actor-critic algorithm returns a policy $\widehat{\pi}$ that satisfies
$\mathrm{SubOpt}(\pi_{\star},\widehat{\pi})=O\Biggl($ $\frac{H^{2}s\log(dNH\delta^{-1})}{\xi\sqrt{N}}$ (13)
$+H^{2}\sqrt{\frac{\log(|\mathcal{A}|)}{N}}+\frac{H^{2}s\sqrt{\epsilon}}{\xi}\Biggl),$
@@ 5.2 Sparse Robust AC with Single Concentrability
$\alpha_{h}=O\Biggl((H-h)\Biggl(\frac{s^{\nicefrac{{1}}{{4}}}\sqrt{\log(dHN/\delta)}}{N^{\nicefrac{{1}}{{4}}}}+\sqrt{\lambda}+\sqrt{\epsilon}\Biggl)\Biggl).$
@@
$\mathrm{SubOpt}(\pi_{\star},\widehat{\pi})=O\Biggl($ $\frac{H^{2}\sqrt{\kappa}s^{\nicefrac{{3}}{{4}}}\sqrt{\log(dHN\delta^{-1})}}{N^{\tfrac{1}{4}}}$
$+H^{2}\sqrt{\frac{\log(|\mathcal{A}|)}{N}}+H^{2}\sqrt{\kappa s\epsilon}\Biggl),$
@@
$\mathrm{SubOpt}(\pi_{\star},\widehat{\pi})=\widetilde{O}(H^{2}\sqrt{\kappa s\epsilon}).$ (15)
@@
$\alpha_{h}=O\Biggl((H-h)\Biggl(\frac{\sqrt{\log(dHN/\delta)}}{N^{\nicefrac{{1}}{{4}}}}+\epsilon^{\nicefrac{{1}}{{4}}}+\sqrt{\lambda}+\sqrt{\epsilon}\Biggl)\Biggl).$
@@
$\mathrm{SubOpt}(\pi_{\star},\widehat{\pi})=O\Biggl($ $\frac{H^{2}\sqrt{\kappa s}\sqrt{\log(dHN\delta^{-1})}}{N^{\tfrac{1}{4}}}$ (16)
$+H^{2}\sqrt{\frac{\log(|\mathcal{A}|)}{N}}+H^{2}\sqrt{\kappa s}\epsilon^{\frac{1}{4}}\Biggl),$
@@
$\mathrm{SubOpt}(\pi_{\star},\widehat{\pi})=\widetilde{O}(H^{2}\sqrt{\kappa s}\epsilon^{\frac{1}{4}}).$ (17)
@@ Appendix D Proof of Section 4
First, by Claim B.1, for any $Q_{h+1}\in\mathcal{Q}_{h+1}$, $\mathcal{P}_{h}^{\pi}(Q_{h+1})$ is $s$-sparse, and $\|\mathcal{P}_{h}^{\pi}(Q_{h+1})\|_{1}\leq(H-h+1)$. Moreover, by Claim B.2, for any trajectory $\tau$, we have $\left|R_{h}+\max_{a^{\prime}\in\mathcal{A}}Q_{h+1}(x^{\prime},a^{\prime})\right|\leq H$. Therefore, using the estimation error in Proposition 3.1 with $\|w\|_{1}\leq 2Hs$ and $\sigma\leq H$, we obtain
$\|\mathcal{R}^{*}_{h}(\underline{Q}_{h+1})-\mathcal{P}^{*}_{h}(\underline{Q}_{h+1})\|_{1}=O\left(\frac{Hs\log(d/\delta)}{\xi\sqrt{N}}+\frac{Hs\sqrt{\epsilon}}{\xi}\right).$ (37)
@@
$=O\left(\frac{H^{2}s\log(d/\delta)}{\xi\sqrt{N}}+\frac{H^{2}s\sqrt{\epsilon}}{\xi}\right).$
@@ Appendix E Proof of Section 5
$\beta_{h}=O\!\left(\frac{(H-h)s\log(dHN\delta^{-1})}{\xi\sqrt{N}}+\frac{(H-h)s\sqrt{\epsilon}}{\xi}\right).$
@@
Putting things together. Let $\widehat{\pi}$ be the mixture of policies $\{\pi_{1},...,\pi_{T}\}$ and $T=N$, we have
$V_{1}^{\pi_{\star}}(x_{1})-V_{1}^{\pi_{t}}(x_{1})\leq 2\sum_{h=1}^{H}\beta_{h}+4H^{2}\sqrt{\frac{\log\mathcal{A}}{N}}$ (47)
@@
$V_{1}^{\pi_{\star}}(x_{1})-V_{1}^{\pi_{t}}(x_{1})=O\!\left(\frac{H^{2}s\log(dNH\delta^{-1})}{\xi\sqrt{N}}+H^{2}\sqrt{\frac{\log(|\mathcal{A}|)}{N}}+\frac{H^{2}s\sqrt{\epsilon}}{\xi}\right),$ (48)
@@
$\alpha_{h}^{2}=O\!\left(\frac{\sqrt{s}(H-h)^{2}\log(dH\delta^{-1})}{\sqrt{N}}+(H-h)^{2}\epsilon+(H-h)^{2}(\lambda+\epsilon)\right),$
@@
${\alpha^{\prime}}_{h}^{2}=O\!\left(\frac{(H-h)^{2}\log(dH\delta^{-1})}{\sqrt{N}}+(H-h)^{2}\sqrt{\epsilon}+(H-h)^{2}(\lambda+\epsilon)\right).$
@@
By Proposition 3.2, for all $h\in[H]$, the SRLE2 oracle ensures that
$\|\mathcal{E}_{h}^{\pi}(Q_{h+1})\|_{\Sigma_{h}}^{2}=O\!\left(\frac{(H-h)^{2}\sqrt{s}\log(dH/\delta)}{\sqrt{N}}+(H-h)^{2}\epsilon\right),$
@@
$\|\mathcal{E}_{h}^{\pi}(Q_{h+1})\|_{\widehat{\Sigma}_{h}}^{2}=O\!\left(\|\mathcal{E}_{h}^{\pi}(Q_{h+1})\|_{\Sigma_{h}}^{2}+(H-h)^{2}(\epsilon+\lambda)\right).$ (51)
@@
Let $e=\mathcal{E}_{h}^{\pi}(Q_{h+1})$. We have $\|e\|_{1}=O(H-h)$ and $\|e\|_{0}\leq 2s$. Let $S$ denote the sparsity support of $e$. Then
@@
$=e^{\top}_{S}\!\left[\frac{1}{N}\sum_{i=1}^{N}\phi_{i}\phi_{i}^{\top}+(\lambda+\epsilon)I\right]e_{S}$
$=e^{\top}_{S}\!\left[\underbrace{\frac{1}{N}\sum_{i=1}^{(1-\epsilon)N}\widetilde{\phi}_{i}\widetilde{\phi}_{i}^{\top}}_{(1-\epsilon)\widetilde{\Sigma}_{h},\text{ clean data}}+\underbrace{\frac{1}{N}\sum_{i=1}^{\epsilon N}\phi_{i}^{\prime}\phi_{i}^{\prime\top}}_{\text{corrupted data}}+(\lambda+\epsilon)I\right]e_{S}.$
@@
$\frac{1}{N}\sum_{i=1}^{\epsilon N}e^{\top}\phi_{i}^{\prime}\phi_{i}^{\prime\top}e\leq\frac{1}{N}\sum_{i=1}^{\epsilon N}(\|\phi_{i}^{\prime}\|_{\infty}\|e\|_{1})^{2}\leq\epsilon\|e\|_{1}^{2}\leq O((H-h)^{2}\epsilon).$
@@
$\left[\underbrace{\frac{1}{N}\sum_{i=1}^{(1-\epsilon)N}\widetilde{\phi}_{i}\widetilde{\phi}_{i}^{\top}}_{(1-\epsilon)\widetilde{\Sigma}_{h},\text{ clean data}}+(\lambda+\epsilon)I\right]_{S}=O([\Sigma_{h}+(\lambda+\epsilon)I]_{S}).$
@@
$\|e\|_{\widehat{\Sigma}_{h}}^{2}=O(\|e\|_{\Sigma_{h}}^{2}+(H-h)^{2}(\lambda+\epsilon)).$
@@
$\frac{1}{T}\sum_{t=1}^{T}\!\left[V_{1,M_{t}}^{\pi_{\star}}(x_{1})-V_{1,M_{t}}^{\pi_{t}}(x_{1})\right]\leq 4H\sqrt{\frac{\log\mathcal{A}}{T}}$ (60)
Therefore, let $\widehat{\pi}$ be the mixture of policies $\{\pi_{1},...,\pi_{T}\}$. Let $T=N$. We have that
$V_{1}^{\pi_{\star}}(x_{1})-V_{1}^{\widehat{\pi}}(x_{1})\leq 2\sqrt{6\kappa}\sum_{h=1}^{H}\alpha_{h}+4H\sqrt{\frac{\log\mathcal{A}}{N}},$ (61)

### 12 May 2026 v3

@@ 1.1 Related works
(Ye et al., 2023) $\widetilde{O}\!\left(\tfrac{H^{3}\sqrt{d}\epsilon}{\xi}\right)$ $\mathrm{poly}(d,H)$ Uniform coverage
(Zhang et al., 2021) $\widetilde{O}\!\left(\tfrac{H^{3}\sqrt{d}\epsilon}{\xi}\right)$ $\mathrm{poly}(d,H)$ Uniform coverage
Ours $\widetilde{O}\!\left(\tfrac{H^{3}s\sqrt{\epsilon}}{\xi}\right)$ $\mathrm{poly}(s,H)$ Uniform coverage
(Zhang et al., 2021) $\widetilde{O}\!\left(H^{3}\sqrt{d\kappa\epsilon}\right)$ $\mathrm{poly}(d,H)$ Single-policy concentrability
Ours $\widetilde{O}\!\left(H^{3}\sqrt{s\kappa\epsilon}\right)$ $\mathrm{poly}(s,H)$ Single-policy concentrability
@@
In this paper, we consider data poisoning attacks in offline RL (Zhang et al., 2021). Previous work on corruption-robust offline RL has primarily considered the low-dimensional setting where $N\gg d$ (Zhang et al., 2021; Ye et al., 2023; Mandal et al., 2025). In contrast, this work focuses on the high-dimensional setting $d>N$. In particular, we employ a value function estimation approach (Zhang et al., 2021) based on sparse robust estimation (Merad and Gaïffas, 2022). We find that, under single-policy concentrability, in sparse settings, integrating sparse oracles into LSVI-type algorithms is considerably more challenging and may yield suboptimal bounds that depend polynomially on $d$, in contrast to prior work in the low-dimensional setting (Zhang et al., 2021). Moreover, we demonstrate that sparse robust estimation can be naturally integrated into the AC framework, effectively eliminating polynomial dependence on $d$. A systematic comparison of our results with previous work on corruption-robust offline RL is provided in Table 1.
@@ 2 PRELIMINARIES
We consider the offline setting with contaminated data. Let $\widetilde{\mathcal{D}}$ be a clean dataset consisting of $N$ trajectories, i.e., $\widetilde{\mathcal{D}}=\{(x^{\tau}_{h},a_{h}^{\tau},R_{h}^{\tau})\}_{\tau=1,h=1}^{N,H}$. Suppose there is an adversary who observes the dataset $\widetilde{\mathcal{D}}$ and may arbitrarily corrupt up to $\epsilon N$ trajectories, resulting in a corrupted dataset $\mathcal{D}$. We split the dataset into $H$ disjoint subsets $\mathcal{D}_{1},\ldots,\mathcal{D}_{H}$, and use $\mathcal{D}_{h}$ only for the stage-$h$ regression. In this work, regarding clean data set, we assume there is an exploratory policy $\pi_{\nu}$, with the occupancy probability $\nu=(\nu_{h})_{h=1}^{H}$. Define $\Sigma_{h}\triangleq\mathbb{E}_{\nu_{h}}\!\left[\phi(x,a)\phi(x,a)^{\top}\right]$ as the covariance matrix of the underlying data distribution $\nu$ at horizon $h$. We introduce an assumption on the distribution $\nu$, which is crucial for our main result.
@@ 3.3 Using SRLE Estimators in Offline Reinforcement Learning
$\mathcal{D}^{\pi}_{h}\triangleq\!\left\{\!\left(\underbrace{\phi(x_{h}^{\tau},a_{h}^{\tau})}_{z_{i}},\;\\ \underbrace{R_{h}^{\tau}+\mathbb{E}_{\pi_{h+1}}\!\left[F(x^{\tau}_{h+1},a)\right]}_{y_{i}}\right)\right\}_{\tau=1}^{N/H}.$
@@
$\mathcal{D}^{*}_{h}$ $\triangleq\!\left\{\!\left(\phi(x_{h}^{\tau},a_{h}^{\tau}),\;R_{h}^{\tau}+\max_{a\in\mathcal{A}}F(x^{\tau}_{h+1},a)\right)\right\}_{\tau=1}^{N/H}.$ (6)
@@ 4.1 Sparse Robust LSVI with Uniform Coverage
$\mathrm{SubOpt}(\pi_{\star},\widehat{\pi})=O\left(\frac{H^{3}s\log(d/\delta)}{\xi\sqrt{N}}+\frac{H^{3}s\sqrt{\epsilon}}{\xi}\right).$ (7)
@@ 5.1 Sparse Robust AC with Uniform Coverage
Suppose Assumption 2.1 and 2.2 hold. Then, with step size $\eta=\sqrt{\frac{\log\mathcal{A}}{N}}$, after $T=N/H$ iterations the actor-critic algorithm returns a policy $\widehat{\pi}$ that satisfies
$\mathrm{SubOpt}(\pi_{\star},\widehat{\pi})=O\Biggl($ $\frac{H^{3}s\log(dNH\delta^{-1})}{\xi\sqrt{N}}$ (13)
$+H^{3}\sqrt{\frac{\log(|\mathcal{A}|)}{N}}+\frac{H^{3}s\sqrt{\epsilon}}{\xi}\Biggl),$
@@ 5.2 Sparse Robust AC with Single Concentrability
$\alpha_{h}^{2}=O\Biggl(\frac{\sqrt{s}H^{3}\log(dH\delta^{-1})}{\sqrt{N}}+H^{3}\epsilon+H^{3}(\lambda+\epsilon)\Biggl).$
@@
$\mathrm{SubOpt}(\pi_{\star},\widehat{\pi})=O\Biggl($ $\frac{H^{3}\sqrt{\kappa}s^{\nicefrac{{3}}{{4}}}\sqrt{\log(dHN\delta^{-1})}}{N^{\tfrac{1}{4}}}$
$+H^{3}\sqrt{\frac{\log(|\mathcal{A}|)}{N}}+H^{3}\sqrt{\kappa s\epsilon}\Biggl),$
@@
$\mathrm{SubOpt}(\pi_{\star},\widehat{\pi})=\widetilde{O}(H^{3}\sqrt{\kappa s\epsilon}).$ (15)
@@
$\alpha_{h}=O\!\left(\frac{H^{3}\log(dH\delta^{-1})}{\sqrt{N}}+H^{3}\sqrt{\epsilon}+H^{3}(\lambda+\epsilon)\right).$
@@
$\mathrm{SubOpt}(\pi_{\star},\widehat{\pi})=O\Biggl($ $\frac{H^{3}\sqrt{\kappa s}\sqrt{\log(dHN\delta^{-1})}}{N^{\tfrac{1}{4}}}$ (16)
$+H^{3}\sqrt{\frac{\log(|\mathcal{A}|)}{N}}+H^{3}\sqrt{\kappa s}\epsilon^{\frac{1}{4}}\Biggl),$
@@
$\mathrm{SubOpt}(\pi_{\star},\widehat{\pi})=\widetilde{O}(H^{3}\sqrt{\kappa s}\epsilon^{\frac{1}{4}}).$ (17)
@@ 6 CONCLUSION
It would also be interesting to study multi-agent Markov games under sparsity. The structure of sparse linear MDPs can be naturally extended to multi-agent Markov games. However, existing algorithms for solving offline Markov games are based on value iteration based methods, and handling sparsity will require generalization of our actor-critic framework to multi-agent setting. Furthermore, recent work Nika et al. (2024) on corruption robustness in offline Markov games has highlighted that data coverage assumptions play an important role on achievable error rates, and it would be interesting to explore how sparsity influences such data coverage assumptions.
Acknowledgements
The work of Andi Nika and Goran Radanovic was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – project number 467367360.
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes]
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes]
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Not Applicable]
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes]
    2. (b)
      Complete proofs of all theoretical results. [Yes]
    3. (c)
      Clear explanations of any assumptions. [Yes]
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Not Applicable]
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Not Applicable]
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Not Applicable]
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Not Applicable]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Not Applicable]
    2. (b)
      The license information of the assets, if applicable. [Not Applicable]
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Not Applicable]
    4. (d)
      Information about consent from data providers/curators. [Not Applicable]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
Sparse Offline Reinforcement Learning with Corruption Robustness
Sparse Offline Reinforcement Learning with Corruption Robustness
@@ Appendix D Proof of Section 4
First, by Claim B.1, for any $Q_{h+1}\in\mathcal{Q}_{h+1}$, $\mathcal{P}_{h}^{\pi}(Q_{h+1})$ is $s$-sparse, and $\|\mathcal{P}_{h}^{\pi}(Q_{h+1})\|_{1}\leq(H-h+1)$. Moreover, by Claim B.2, for any trajectory $\tau$, we have $\left|R_{h}+\max_{a^{\prime}\in\mathcal{A}}Q_{h+1}(x^{\prime},a^{\prime})\right|\leq H$. Note that $|\mathcal{D}_{h}|=N/H$ and the worst-case contamination level in $\mathcal{D}_{h}$ is $H\epsilon$. Therefore, using the estimation error in Proposition 3.1 with $\|w\|_{1}\leq 2Hs$ and $\sigma\leq H$, we obtain
$\|\mathcal{R}^{*}_{h}(\underline{Q}_{h+1})-\mathcal{P}^{*}_{h}(\underline{Q}_{h+1})\|_{1}=O\left(\frac{H^{\tfrac{3}{2}}s\log(d/\delta)}{\xi\sqrt{N}}+\frac{H^{2}s\sqrt{\epsilon}}{\xi}\right).$ (37)
@@
$=O\left(\frac{H^{3}s\log(d/\delta)}{\xi\sqrt{N}}+\frac{H^{3}s\sqrt{\epsilon}}{\xi}\right).$
@@ Appendix E Proof of Section 5
$\beta_{h}=O\!\left(\frac{H^{2}s\log(dHN\delta^{-1})}{\xi\sqrt{N}}+\frac{H^{2}s\sqrt{\epsilon}}{\xi}\right).$
@@
Putting things together. Let $\widehat{\pi}$ be the mixture of policies $\{\pi_{1},...,\pi_{T}\}$ and $T=N/H$, we have
$V_{1}^{\pi_{\star}}(x_{1})-V_{1}^{\pi_{t}}(x_{1})\leq 2\sum_{h=1}^{H}\beta_{h}+4H^{3}\sqrt{\frac{\log\mathcal{A}}{N}}$ (47)
@@
$V_{1}^{\pi_{\star}}(x_{1})-V_{1}^{\pi_{t}}(x_{1})=O\!\left(\frac{H^{3}s\log(dNH\delta^{-1})}{\xi\sqrt{N}}+H^{3}\sqrt{\frac{\log(|\mathcal{A}|)}{N}}+\frac{H^{3}s\sqrt{\epsilon}}{\xi}\right),$ (48)
@@
$\alpha_{h}^{2}=O\!\left(\frac{\sqrt{s}H^{3}\log(dH\delta^{-1})}{\sqrt{N}}+H^{3}\epsilon+H^{3}(\lambda+\epsilon)\right),$
@@
${\alpha^{\prime}}_{h}^{2}=O\!\left(\frac{H^{3}\log(dH\delta^{-1})}{\sqrt{N}}+H^{3}\sqrt{\epsilon}+H^{3}(\lambda+\epsilon)\right).$
@@
Note that dual to data spliting, the worst case contamination level in set $\mathcal{D}_{h}$ is $H\epsilon$. By Proposition 3.2, for all $h\in[H]$, the SRLE2 oracle ensures that
$\|\mathcal{E}_{h}^{\pi}(Q_{h+1})\|_{\Sigma_{h}}^{2}=O\!\left(\frac{H^{3}\sqrt{s}\log(dH/\delta)}{\sqrt{N}}+H^{3}\epsilon\right),$
@@
$\|\mathcal{E}_{h}^{\pi}(Q_{h+1})\|_{\widehat{\Sigma}_{h}}^{2}=O\!\left(\|\mathcal{E}_{h}^{\pi}(Q_{h+1})\|_{\Sigma_{h}}^{2}+H^{3}(\epsilon+\lambda)\right).$ (51)
@@
Let $e=\mathcal{E}_{h}^{\pi}(Q_{h+1})$. Note that, due to data splitting, $Q_{h+1}$ does not dependent on data set $\mathcal{D}_{h}$. We have $\|e\|_{1}=O(H-h)$ and $\|e\|_{0}\leq 2s$. Let $S$ denote the sparsity support of $e$. Then
@@
$=e^{\top}_{S}\!\left[\frac{H}{N}\sum_{i=1}^{N/H}\phi_{i}\phi_{i}^{\top}+(\lambda+\epsilon)I\right]e_{S}$
$=e^{\top}_{S}\!\left[\underbrace{\frac{H}{N}\sum_{i=1}^{(1-\epsilon)N/H}\widetilde{\phi}_{i}\widetilde{\phi}_{i}^{\top}}_{(1-\epsilon)\widetilde{\Sigma}_{h},\text{ clean data}}+\underbrace{\frac{H}{N}\sum_{i=1}^{\epsilon N/H}\phi_{i}^{\prime}\phi_{i}^{\prime\top}}_{\text{corrupted data}}+(\lambda+\epsilon)I\right]e_{S}.$
@@
$\frac{H}{N}\sum_{i=1}^{\epsilon N/H}e^{\top}\phi_{i}^{\prime}\phi_{i}^{\prime\top}e\leq\frac{H}{N}\sum_{i=1}^{\epsilon N/H}(\|\phi_{i}^{\prime}\|_{\infty}\|e\|_{1})^{2}\leq\epsilon\|e\|_{1}^{2}\leq O((H-h)^{2}\epsilon).$
@@
$\left[\underbrace{\frac{H}{N}\sum_{i=1}^{(1-\epsilon)N/H}\widetilde{\phi}_{i}\widetilde{\phi}_{i}^{\top}}_{(1-\epsilon)\widetilde{\Sigma}_{h},\text{ clean data}}+(\lambda+\epsilon)I\right]_{S}=O([\Sigma_{h}+(\lambda+H\epsilon)I]_{S}).$
@@
$\|e\|_{\widehat{\Sigma}_{h}}^{2}=O(\|e\|_{\Sigma_{h}}^{2}+H^{3}(\lambda+\epsilon)).$
@@
$\frac{1}{T}\sum_{t=1}^{T}\!\left[V_{1,M_{t}}^{\pi_{\star}}(x_{1})-V_{1,M_{t}}^{\pi_{t}}(x_{1})\right]\leq 4H^{2}\sqrt{\frac{\log|\mathcal{A}|}{T}}$ (60)
Therefore, let $\widehat{\pi}$ be the mixture of policies $\{\pi_{1},...,\pi_{T}\}$. Let $T=N/H$. We have that
$V_{1}^{\pi_{\star}}(x_{1})-V_{1}^{\widehat{\pi}}(x_{1})\leq 2\sqrt{6\kappa}\sum_{h=1}^{H}\alpha_{h}+4H^{3}\sqrt{\frac{\log|\mathcal{A}|}{N}},$ (61)